### 06 Apr 2026 v1

@@ A.1 Proof of Proposition 1
For any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and any direction $h_{\hat{x}}\in\hat{\mathbb{E}}$, define $h=(h_{\hat{x}},0)\in\mathbb{E}\times\mathbb{E}$ and let $\varrho(t)=D^{2}f(w+th)[h,h]=D^{2}_{\hat{x}\hat{x}}f(w+th)[h_{\hat{x}},h_{\hat{x}}$]. By the $\alpha$-self-concordant convex-concave property of $f$, we have
@@
$D^{3}_{\hat{x}\hat{x}\hat{x}}f(\hat{x},s)[h_{\hat{x}},h_{\hat{x}},h_{\hat{x}}]\leq\frac{2}{\alpha^{1/2}}\left(D^{2}_{\hat{x}\hat{x}}f(\hat{x},s)[h_{\hat{x}},h_{\hat{x}}]\right)^{3/2}.$
@@
$\xi_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\big\langle\nabla_{\hat{x}}f(w),\big(D^{2}_{\hat{x}\hat{x}}f(w)\big)^{-1}\nabla_{\hat{x}}f(w)\big\rangle\Big)^{1/2},$ (145)
@@
$\tilde{\delta}_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\langle d,D^{2}_{\hat{x}\hat{x}}f(\hat{x},s)\,d\rangle\Big)^{1/2}.$ (146)
@@
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-(1-3\xi(w))^{1/3}.$ (149)
Further, if $\xi(w)\leq 0.1$, then
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-0.7^{1/3}<0.2,$ (150)

### 21 May 2026 v2

@@ A.1 Proof of Proposition 1
For any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and any direction $h_{\hat{x}}\in\hat{\mathbb{E}}$, define $h=(h_{\hat{x}},0)\in\hat{\mathbb{E}}\times\mathbb{E}$ and let $\varrho(t)=D^{2}f(w+th)[h,h]=D^{2}_{\hat{x}\hat{x}}f(w+th)[h_{\hat{x}},h_{\hat{x}}$]. By the $\alpha$-self-concordant convex-concave property of $f$, we have
@@
$|D^{3}_{\hat{x}\hat{x}\hat{x}}f(w)[h_{\hat{x}},h_{\hat{x}},h_{\hat{x}}]|\leq\frac{2}{\alpha^{1/2}}\left(D^{2}_{\hat{x}\hat{x}}f(w)[h_{\hat{x}},h_{\hat{x}}]\right)^{3/2}.$
Moreover, for any $w\in\hat{\mathbb{E}}\times\mathbb{E}$ and any nonzero direction $h_{\hat{x}}\in\hat{\mathbb{E}}$,
$D^{2}_{\hat{x}\hat{x}}f(w)[h_{\hat{x}},h_{\hat{x}}]>0.$
@@
$\xi_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\big\langle\nabla_{\hat{x}}f(w),\big(D^{2}_{\hat{x}\hat{x}}f(w)\big)^{-1}\nabla_{\hat{x}}f(w)\big\rangle\Big)^{1/2},$
@@
$\tilde{\delta}_{\hat{x}}(w)=\Big(\tfrac{1}{\alpha}\langle d,D^{2}_{\hat{x}\hat{x}}f(w)\,d\rangle\Big)^{1/2}.$
@@
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-(1-3\xi(w))^{1/3}.$
Furthermore, if $\xi(w)\leq 0.1$, then
$\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}\leq 1-0.7^{1/3}<0.2,$